OpenStudy (anonymous):

Add the fractions and write your answer in lowest terms. two-fifths + one-ninth

OpenStudy (zzr0ck3r):

\(\frac{2}{5}+\frac{1}{9}=\frac{18}{45}+\frac{5}{45}\)

OpenStudy (zzr0ck3r):

Do you see how I got the common denominator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yeah sorry

OpenStudy (anonymous):

\[\frac{ 23}{ 45 }\]